Use the Laws of logarithms to write the following as a logarithm of a single expression. When writing log(x) , you need to use parentheses and the expression should be typed as log(x).
4log(x)−17log(y) =

Answer by Theo(13342)

4log(x) - 17log(y) = log(x^4) - log(y^17) = log(x^4/y^17)

concepts used:

alog(x) = log(x^a)

log(x) - log(y) = log(x/y)